Consider the following integral:


I =  ∫ 1 / (√2x2-x2) dx


(a) complete the square of f(x) = 2x2-x2

(b) use (a) together with the method of trigonometric substitution to determine the integral.


1
Expert's answer
2021-08-08T17:57:19-0400

"\displaystyle
(a)\\ \textsf{Completing the square}\\
2x - x^2 = -(-2x + x^2) = -(1 - 2x + x^2) + 1 = -(x - 1)^2 + 1\\

(b)\\ \textsf{Under the substitution}\,\, x - 1 = \sin\theta\\
\begin{aligned}
\int \frac{1}{\sqrt{2x - x^2}}\mathrm{d}x &= \int \frac{1}{\sqrt{1 - (x - 1)^2}}\mathrm{d}x
\\&= \int \frac{1}{\sqrt{1 - \sin^2{\theta}}}\cos{\theta}\mathrm{d}\theta
\\&= \int \frac{1}{\sqrt{\cos^2{\theta}}}\cos{\theta}\mathrm{d}\theta
\\&= \int \mathrm{d}\theta = \theta + C
\\&= \arcsin(x - 1) + C
\end{aligned}"